# What Drives Issue Resolution Speed? An Empirical Study of Scientific Workflow Systems on GitHub

### v1

@@ I Introduction
In today’s research ecosystem, scientific workflows have become a cornerstone for enabling reproducible, scalable, and automated data-intensive experiments [pouchard2019computational]. Scientific workflow systems (SWSs) provide the infrastructure to design, execute, monitor, and share workflows composed of inter-dependent tasks, data transformations, and computational resources [atkinson2017scientific, barker2007scientific]. Examples include platforms such as Galaxy [galaxy2022galaxy], Nextflow [di2017nextflow], Pegasus [deelman2015pegasus], and Snakemake [koster2012snakemake], which support domains from bioinformatics to remote sensing. These systems allow researchers to model complex analytic pipelines as workflows, manage data dependencies and provenance, automate execution on heterogeneous computing resources (e.g., HPC clusters, cloud), and thus shift attention from manual orchestration toward higher-level scientific insight [liew2016scientific]. However, like any complex software system, SWSs require continuous maintenance. Developers need to fix bugs, implement new features, assist users, update dependencies, and adapt workflows as technologies evolve. Recent studies also highlight that maintaining SWSs comes with unique challenges, including managing complex dependencies, data structures, execution environments, error handling, and ensuring reproducibility [alam2025empirical].
@@
Given these demands, one important aspect of SWS quality and sustainability is how efficiently issues reported by users or contributors are managed and resolved. In open‐source ecosystems such as GitHub, issue trackers provide a mechanism through which bug reports, enhancement requests, questions, and other maintenance tasks are communicated, triaged, addressed, and closed. Prior studies on general software projects have shown that factors such as contributor activity, workload distribution, labeling practices, and the use of structured templates can significantly affect how quickly and effectively issues are resolved [jarczyk2014github, yang2023users, bijlsma2012faster, kuramoto2024understanding]. Faster issue resolution is often associated with higher user satisfaction, stronger community trust, and greater project adoption. However, SWS projects differ from conventional software in several ways; they are often interdisciplinary, serve research-driven purposes, and must manage complex aspects such as data dependencies, execution environments, and workflow reproducibility [alam2025empirical, bharathi2008characterization]. These distinctive characteristics may shape how issues emerge, are prioritized, and are ultimately resolved. Yet, systematic evidence on what drives issue resolution speed in SWSs remains limited. Understanding these factors is crucial for improving the responsiveness, reliability, and long-term sustainability of SWSs.
@@
  • RQ1: How are the issues managed and addressed in scientific workflow systems? This question examines the overall issue management process to characterize the practices adopted by maintainers and contributors to handle bug reports, enhancement requests, and other issue types.
  • RQ2: What are the relationships between different features and the closure of issues in scientific workflow systems? This research question examines how diverse issue- and contributor-related features influence whether and how quickly issues are resolved in SWSs.
Our study analyzes 21,116 issues from 197 repositories related to most popular SWSs (e.g., Galaxy, Nextflow) to investigate what drives issue resolution speed on GitHub. Overall, 68.91% of issues are closed, with half resolved within 18.09 days, though a long tail of unresolved cases indicates persistent maintenance backlogs. Repositories with more contributors do not necessarily resolve issues faster, while those with more total issues tend to take longer to close them. Nextflow shows the best closure rate (86.27%) and shortest average resolution time (136.42 days), likely due to its centralized governance, whereas Galaxy’s distributed structure led to slower responses. Labeling and assignment practices are key management factors; repositories that use them tend to close issues more efficiently. Statistical analysis confirms that while labeling and assignment have small but significant effects on issue closure, textual attributes and contributor counts have negligible influence. These findings highlight that effective triage, coordination, and governance are the main drivers of timely issue resolution in SWSs.
@@ II-A Scientific Workflow System (SWS)
A Scientific Workflow System (SWS) is an efficient software framework designed to define, execute, and manage complex workflows and associated datasets across diverse computing environments [liu2015survey]. Several well-established SWSs, such as Galaxy [galaxy2022galaxy], Pegasus [deelman2005pegasus], Swift [zhao2007swift], Kepler [altintas2004kepler], Taverna [oinn2004taverna], Snakemake [koster2012snakemake], Chiron [ogasawara2013chiron], and Nextflow [di2017nextflow], are now widely adopted across diverse research domains, including bioinformatics and computational biology, medical and health informatics, data science and machine learning, and computational engineering.
@@ II-B Issue Resolution in Open Source Software (OSS)
Prior research [bijlsma2012faster, pfahl2016improving] has shown that issue resolution speed is a critical indicator of project health and developer responsiveness. Studies on OSS repositories have examined how factors such as issue type, reporter expertise, assignee workload, project popularity, and communication activity affect resolution times [bettenburg2008makes, yu2015wait, nguyen2011impact, mockus2002two]. Other works highlight that social and organizational dynamics, like contributor turnover, governance structure, and project maturity, also shape maintenance outcomes [crowston2008free, vasilescu2015gender]. Several studies have leveraged GitHub data (e.g., issues, pull requests, commits) to analyze collaboration patterns and productivity. For instance, Gousios et al. [gousios2014exploratory] and Tsay et al. [tsay2014influence] studied pull request acceptance and developer interaction, while Zhou et al. [zhou2016inflow] investigated temporal patterns in issue handling. These studies emphasize that issue resolution is a multifaceted process influenced by both technical (e.g., code complexity, dependencies) and social (e.g., communication, community size) factors. However, despite the growing importance, SWSs remain largely unexplored in this context, and empirical evidence on what drives issue resolution speed in SWSs is still lacking.
@@ II-C Research Gap
Compared to general-purpose OSS, SWS projects exhibit distinct characteristics, more specialized developer communities, domain-specific user bases, and a strong emphasis on reproducibility and validation [hannay2009scientists, katz2021taking]. SWSs, in particular, must balance usability for domain scientists with technical robustness, which often introduces unique challenges in software maintenance and evolution [alam2025empirical]. However, there is a lack of empirical studies investigating issue management practices and developer responsiveness within this class of projects.
While prior work [nguyen2011impact, ossher2010automated] has explored issue resolution dynamics in general OSS, there is a lack of understanding of how these dynamics manifest in SWSs. This study fills that gap by analyzing issue resolution speed across SWS projects, identifying the factors that drive faster resolution.
@@ III Study Design
To examine issue resolution dynamics within SWSs, we focus on three widely adopted platforms: Galaxy [^1] https://github.com/galaxyproject, Nextflow [^2] https://github.com/nextflow-io, and Snakemake [^3] https://github.com/snakemake. These systems are representative of the broader SWS landscape, each supported by a large and active developer community and exhibiting extensive issue-tracking activity on GitHub [alam2025empirical]. They also embody distinct paradigms of workflow design and execution: Galaxy provides a graphical, web-based environment that simplifies workflow construction for domain scientists [galaxy2022galaxy]; Nextflow introduces a domain-specific language for portable and scalable workflow execution across heterogeneous environments [di2017nextflow]; and Snakemake offers a Python-based framework emphasizing automation and reproducibility [koster2012snakemake]. The diversity of these systems enables a comprehensive examination of issue management practices across varying design philosophies and user communities. Moreover, their open-source nature and long development histories (spanning over a decade) provide transparent, longitudinal datasets suitable for our analysis.
At the time of data collection (September 2025), the Galaxy project supported 147 repositories, Nextflow had 106 repositories, and Snakemake had 63 repositories, resulting in a total of 316 repositories. Among 316 repositories, we find that 197 of them contain issues. To answer our research questions, we collect issue data from these 197 repositories using the GitHub API. In total, we retrieve 21,116 issues. For each issue, we capture all available fields, including the issue’s title, body, creation and closing timestamps, state (open or closed), labels, and discussion comments. In addition to issue-level information, we collect associated repository metadata, such as the number of stars, forks, and contributors, to characterize project popularity and community engagement.
For RQ1, we aim to understand how maintainers and contributors handle issues. This includes examining issue types, labeling and assignment practices, discussion activity, and closure trends across repositories. The goal is to characterize maintenance workflows and identify patterns in issue management and collaboration practices. For RQ2, we investigate how issue- and contributor-level attributes, such as the presence of labels, assignees, and discussions, relate to whether and how quickly issues are resolved. We conduct statistical comparisons between open and closed issues and use correlation-based modeling to identify features most strongly associated with faster resolution. The detailed analytical procedures and results for each research question are presented in the section IV.
Replication Package: The dataset and scripts necessary to replicate this study are available the replication package [random_2025_17555352].
@@ IV-A RQ1: How are issues addressed and managed?
Motivation: Effective issue management is vital for ensuring the quality, usability, and long-term sustainability of SWSs. In GitHub, issues serve as the primary communication channel through which users and developers report bugs, request features, and coordinate maintenance activities. The way these issues are labeled, discussed, assigned, and resolved reflects both the technical maturity of a project and the health of its contributing community. SWSs bring together interdisciplinary contributors, ranging from software engineers to domain scientists, who often differ in their technical expertise, collaboration styles, and expectations. Alam et al. [alam2025empirical] observe that users of SWSs raise issues for diverse reasons, including bug fixing, dependency management, documentation improvements, system redesign, and API migration. While some issues can be easily resolved by users (e.g., package installation or configuration problems), others (e.g., system redesign and API migration) demand careful coordination and long-term planning from maintainers. Neglecting these issues can hinder users from effectively applying SWSs in their scientific analyses. Given the rapid expansion of the SWS landscape, understanding how issues are managed and addressed is essential for improving maintainability and supporting sustainable community practices.
How are the issues addressed? We aim to examine how issues are managed and addressed in SWSs. These issues are divided into open and closed categories. Open issues represent ongoing or unresolved problems, whereas closed issues indicate those that have been resolved. Among all analyzed repositories, 197 contain issues, and these form the basis of our subsequent analysis. Using the GitHub API, we collect a total of 21,166 issues from the selected repositories. Of these, 6,581 (31.09%) are open and 14,585 (68.91%) are closed. On average, each repository contains 107.44 issues. Notably, 36 repositories (18.27%) have closed all their issues. Furthermore, 44 repositories (22.34%) exhibit a closed issue rate (CIR) exceeding 90%, with an average of 17.25 issues each. This indicates that repositories with a CIR greater than 0.9 generally have a smaller number of total issues compared to others. The results of the closed issue rate are presented in Figure 1(a).
@@
RQ1 Summary In our dataset, 68.91% of issues are closed, with a median resolution time of 18.09 days. At the repository level, 40.61% of repositories use labels to categorize issues, and 50.76% assign issues to contributors. At the issue level, 54.09% of issues carry one or more labels, while 20.39% have at least one assignee. Notably, among the top ten most frequently used labels, only three correspond to GitHub’s default set, while seven are community-defined (e.g., workflows, paper cuts, backend)
IV-B RQ2: What are the relationships between different features and the closure of issues in scientific workflow systems?
Motivation: This question aims to explore the relationships between various features of issues and their closure in SWS repositories. SWSs are developed within interdisciplinary teams that include domain scientists and software engineers, leading to diverse development practices and maintenance challenges. Understanding how factors such as issue characteristics, contributor activity, and project attributes relate to issue closure can shed light on what drives efficient problem resolution in these projects. The answer to this research question can provide practical suggestions for both repository maintainers and issue raisers to improve the management, prioritization, and resolution of issues in SWS repositories.
@@
  • has-code: Identifies whether the issue body includes code blocks. Code snippets, such as reproducing examples or error traces can help maintainers understand and reproduce the reported problem, aiding diagnosis and resolution.
@@
  • num-contributors: Represents the total number of contributors in a repository. Repositories with more active contributors are intuitively more capable of addressing issues efficiently.
Feature Measurement and Statistical Test. We measure the length of each issue title and body by applying the built-in Python len() function. Using regular expressions, we further detect whether the issue description contained code blocks or URLs. All issues are then grouped into two categories, open and closed. For each issue, we extract the six structural and management-related features, including has-label, has-assignee, title-length, and body-length. The number of contributors in a repository is obtained by querying the GitHub API to fetch metadata about the repository. Finally, to determine whether these feature distributions significantly differed between open and closed issues, we apply the non-parametric Wilcoxon rank-sum test [wilcoxon1945individual].
TABLE I: The significance of the differences between the feature values of open and closed issues (Wilcoxon rank-sum test and Cohen’s effect sizes $\delta$).
@@
In addition to statistical significance testing, we also compute Cohen’s effect size ($\delta$) to quantify the magnitude of differences between open and closed issues. While p-values indicate whether the observed differences are statistically significant, effect sizes capture their practical impact, that is, how meaningful these differences are in real-world terms. Following established conventions, values of $|\delta|<0.2$ are considered negligible, $0.2\leq|\delta|<0.5$ as small, $0.5\leq|\delta|<0.8$ as medium, and $|\delta|\geq 0.8$ as large.
Table I summarizes the statistical comparison between open and closed issues across all investigated features. All features show statistically significant differences ($p>0.01$), indicating that each characteristic varies meaningfully between open and closed issues. However, the effect size values ($\delta$) reveal that the magnitude of these differences is generally small or negligible, suggesting that while the differences are statistically detectable, their practical impact is limited.
However, the presence of labels and assigned developers (i.e., has-label and has-assignee) show small but meaningful effects ($\delta$ = 0.20 and 0.23, respectively). This finding aligns with expectations that labeled and assigned issues are more likely to attract attention and thus reach closure. These two features appear to play a modest but tangible role in issue management effectiveness. In contrast, textual and contextual attributes, such as title-length, body-length, has-code, and has-URLs exhibit negligible effect sizes ($\delta$ << 0.2), implying that the amount of written content or inclusion of supporting information (e.g., code snippets, URLs) does not substantially differentiate open from closed issues. Similarly, the number of contributors in a repository shows only a negligible effect, indicating that having more contributors does not necessarily lead to faster or more consistent issue resolution.
@@
RQ2 Summary Our analysis shows that management-related factors, particularly labeling and assignment, have the strongest association with issue closure in SWS repositories, albeit with small effect sizes. In contrast, content-related attributes (e.g., title length, body length, code snippets, URLs) and repository-level characteristics (e.g., number of contributors) show negligible influence. These findings suggest that effective triage and task assignment practices, rather than issue content or team size, play a more meaningful role in driving issue resolution.

### v2

@@ I Introduction
In today’s research ecosystem, scientific workflows have become a cornerstone for enabling reproducible, scalable, and automated data-intensive experiments [27]. Scientific workflow systems (SWSs) provide the infrastructure to design, execute, monitor, and share workflows composed of inter-dependent tasks, data transformations, and computational resources [6, 7]. Examples include platforms such as Galaxy [29], Nextflow [14], Pegasus [13], and Snakemake [19], which support domains from bioinformatics to remote sensing. These systems allow researchers to model complex analytic pipelines as workflows, manage data dependencies and provenance, automate execution on heterogeneous computing resources (e.g., HPC clusters, cloud), and thus shift attention from manual orchestration toward higher-level scientific insight [21]. However, like any complex software system, SWSs require continuous maintenance. Developers need to fix bugs, implement new features, assist users, update dependencies, and adapt workflows as technologies evolve [2]. Recent studies also highlight that maintaining SWSs presents several challenges, including managing complex dependencies, data structures, and execution environments, handling errors, and ensuring reproducibility [1, 4].
@@
Given these demands, one important aspect of SWS quality and sustainability is how efficiently issues reported by users or contributors are managed and resolved. In open‐source ecosystems such as GitHub, issue trackers provide a mechanism through which bug reports, enhancement requests, questions, and other maintenance tasks are communicated, triaged, addressed, and closed. Prior studies on general software projects have shown that factors such as contributor activity, workload distribution, labeling practices, and the use of structured templates can significantly affect how quickly and effectively issues are resolved [17, 32, 10, 20]. Faster issue resolution is often associated with higher user satisfaction, stronger community trust, and greater project adoption.
However, SWS projects differ from conventional software in several ways: SWSs are dataflow-oriented platforms that automatically orchestrate computational research pipelines, fundamentally differing from traditional software in their execution model and user focus. They explicitly represent data dependencies as directed acyclic graphs, enabling automatic parallelization across distributed environments, while traditional software uses imperative programming with obscured data flow. Scientific workflows prioritize reproducibility through automatic provenance tracking [3] and are built for petascale data and long-running jobs across HPC and cloud infrastructure. They target domain scientists via visual interfaces and reusable containerized tasks, whereas traditional software emphasizes business logic and developer-centric design. They are often interdisciplinary, serve research-driven purposes, and must manage complex aspects such as data dependencies, execution environments, and workflow reproducibility [1, 9]. These distinctive characteristics may shape how issues emerge, are prioritized, and are ultimately resolved. Yet, systematic evidence on what drives issue resolution speed in SWSs remains limited. Understanding these factors is crucial for improving the responsiveness, reliability, and long-term sustainability of SWSs.
@@
  • RQ1: How are the issues managed and addressed in scientific workflow systems? This RQ investigates the overall issue management process, aiming to characterize the practices and strategies used by maintainers and contributors in SWSs.
  • RQ2: What are the relationships between different features and the closure of issues in scientific workflow systems? This RQ examines how diverse issue- and contributor-related features influence whether and how quickly issues are resolved in SWSs.
Our study analyzes 21,116 issues from 197 repositories related to popular SWSs (e.g., Galaxy, Nextflow) to investigate what drives issue resolution speed on GitHub. Overall, 68.91% of issues are closed, with half resolved within 18.09 days, though a long tail of unresolved cases indicates persistent maintenance backlogs. Repositories with more contributors do not necessarily resolve issues faster, while those with more total issues tend to take longer to close them. Nextflow shows the best closure rate (86.27%) and shortest average resolution time (136.42 days), likely due to its centralized governance, whereas Galaxy’s distributed structure led to slower responses. Labeling and assignment practices are key management factors; repositories that use them tend to close issues more efficiently. Statistical analysis confirms that while labeling and assignment have small but significant effects on issue closure, textual attributes and contributor counts have negligible influence. These findings highlight that effective triage, coordination, and governance are the main drivers of timely issue resolution in SWSs.
@@ II-A Scientific Workflow System (SWS)
A SWS is an efficient software framework designed to define, execute, and manage complex workflows and associated datasets across diverse computing environments [22]. Several well-established SWSs, such as Galaxy [29], Pegasus [12], Kepler [5], Taverna [25], Snakemake [19], and Nextflow [14], are now widely adopted across diverse research domains, including bioinformatics and computational biology, medical and health informatics, data science and machine learning, and computational engineering.
@@ II-B Issue Resolution in Open Source Software (OSS)
Prior research [10, 26] has shown that issue resolution speed is a critical indicator of project health and developer responsiveness. Studies on OSS repositories have examined how factors such as issue type, reporter expertise, assignee workload, project popularity, and communication activity affect resolution times [8, 33, 24, 23]. Other works highlight that social and organizational dynamics, like contributor turnover, governance structure, and project maturity, also shape maintenance outcomes [11]. Several studies have leveraged GitHub data (e.g., issues, pull requests, commits) to analyze collaboration patterns and productivity. For instance, Gousios et al. [15] and Tsay et al. [30] studied pull request acceptance and developer interaction, while Zhou et al. [34] investigated temporal patterns in issue handling. These studies emphasize that issue resolution is a multifaceted process influenced by both technical (e.g., code complexity, dependencies) and social (e.g., communication, community size) factors. However, despite the growing importance, SWSs remain largely unexplored in this context.
@@ II-C Research Gap
Compared to general-purpose OSS, SWS projects exhibit distinct characteristics, more specialized developer communities, domain-specific user bases, and a strong emphasis on reproducibility and validation [16, 18]. SWSs, in particular, must balance usability for domain scientists with technical robustness, which often introduces unique challenges in software maintenance and evolution [1]. However, there is a lack of empirical studies investigating issue management practices and developer responsiveness within this class of projects. This study addresses this gap by analyzing issue-resolution speed across SWS projects and identifying the factors that drive faster resolution.
@@ III Study Design
To examine issue resolution dynamics within SWSs, we focus on three widely adopted platforms: Galaxy [^1] https://github.com/galaxyproject, Nextflow [^2] https://github.com/nextflow-io, and Snakemake [^3] https://github.com/snakemake. A large and active developer community supports these systems and exhibits extensive issue-tracking activity on GitHub. They also embody distinct paradigms of workflow design and execution: Galaxy provides a graphical, web-based environment that simplifies workflow construction for domain scientists [29]; Nextflow introduces a domain-specific language for portable and scalable workflow execution across heterogeneous environments [14]; and Snakemake offers a Python-based framework emphasizing automation and reproducibility [19]. The diversity of these systems enables a comprehensive examination of issue management practices across varying design philosophies and user communities. Moreover, their open-source nature and long development histories (spanning over a decade) provide transparent, longitudinal datasets suitable for our analysis.
At the time of data collection (September 2025), the Galaxy project supported 147 repositories, Nextflow had 106 repositories, and Snakemake had 63 repositories, resulting in a total of 316 repositories. Among them, 197 contain issues. To answer our RQs, we collect issue data from these 197 repositories using the GitHub API. In total, we retrieved 21,116 issues. For each issue, we capture all available fields, including the issue’s title, body, creation and closing timestamps, state (open or closed), labels, and discussion comments. In addition to issue-level information, we collect associated repository metadata, such as the number of stars, forks, and contributors, to characterize project popularity and community engagement.
For RQ1, we aim to understand how maintainers and contributors handle issues. This includes examining issue types, labeling and assignment practices, discussion activity, and closure trends across repositories. The goal is to characterize maintenance workflows and identify patterns in issue management and collaboration practices. For RQ2, we investigate how issue- and contributor-level attributes, such as the presence of labels, assignees, and discussions, relate to whether and how quickly issues are resolved. We conduct statistical comparisons between open and closed issues and use correlation-based modeling to identify features associated with faster resolution. The detailed analytical procedures and results for each RQ are presented in the section IV.
The replication package is available in online appendix [28].
@@ IV-A RQ1: How are issues addressed and managed?
Motivation: Effective issue management is essential for the quality, usability, and sustainability of SWSs. On GitHub, issues function as the primary medium for reporting bugs, requesting features, and coordinating maintenance, reflecting both project maturity and community health. SWSs involve interdisciplinary contributors with varying expertise, leading to diverse issue types such as bug fixes, dependency management, documentation updates, system redesign, and API migration [1]. While some issues are easily resolved, others require extensive coordination and long-term planning, and neglecting them can hinder the scientific use of SWSs. As the SWS ecosystem continues to grow, understanding issue management practices is critical for improving maintainability and fostering sustainable communities.
How are the issues addressed? We aim to examine how issues are managed and addressed in SWSs. These issues are divided into open and closed categories. Using the GitHub API, we collect a total of 21,166 issues from the selected repositories. Of these, 6,581 (31.09%) are open, and 14,585 (68.91%) are closed. On average, each repository contains 107.44 issues. Notably, 36 repositories (18.27%) have closed all their issues. Furthermore, 44 repositories (22.34%) exhibit a closed issue rate (CIR) exceeding 90%, with an average of 17.25 issues each. This indicates that repositories with a CIR greater than 0.9 generally have a smaller number of total issues compared to others. The results of the closed issue rate are presented in Figure 1(a).
@@
In our dataset, 68.91% of issues are closed, with a median resolution time of 18.09 days. At the repository level, 40.61% of repositories use labels to categorize issues, and 50.76% assign issues to contributors. At the issue level, 54.09% of issues carry one or more labels, while 20.39% have at least one assignee. Notably, among the top ten most frequently used labels, only three correspond to GitHub’s default set, while seven are community-defined (e.g., workflows, paper cuts, backend)
IV-B RQ2:Relationships between different features and the closure of issues
Motivation: This RQ aims to explore the relationships between various features of issues and their closure in SWS repositories. SWSs are developed within interdisciplinary teams that include domain scientists and software engineers, leading to diverse development practices and maintenance challenges. Understanding how factors such as issue characteristics, contributor activity, and project attributes relate to issue closure can shed light on what drives efficient problem resolution in these projects. The answer to this RQ can provide practical suggestions for both repository maintainers and issue raisers to improve the management, prioritization, and resolution of issues in SWS repositories.
@@ IV-B RQ2:Relationships between different features and the closure of issues
  • has-code: Identifies whether the issue body includes code blocks. Code snippets, such as reproducing examples or error traces can help maintainers understand and reproduce the reported problem, aiding resolution.
@@
  • num-contributors: Represents the total number of contributors in a repository; repositories with more active contributors are generally better capable of addressing issues efficiently.
Feature Measurement and Statistical Test. We measure the length of each issue title and body by applying the built-in Python len() function. Using regular expressions, we further detect whether the issue description contained code blocks or URLs. For each issue, we extract the six structural and management-related features, including has-label, has-assignee, title-length, and body-length. The number of contributors in a repository is obtained by querying the GitHub API to fetch metadata about the repository. Finally, to determine whether these feature distributions significantly differed between open and closed issues, we apply the non-parametric Wilcoxon rank-sum test [31].
TABLE I: The significance of the differences between the feature values of open and closed issues
@@
In addition to statistical significance testing, we also compute Cohen’s effect size ($d$) to quantify the magnitude of differences between open and closed issues. While p-values indicate whether the observed differences are statistically significant, effect sizes capture their practical impact, that is, how meaningful these differences are in real-world terms. Following established conventions, values of $|d|<0.2$ are considered negligible, $0.2\leq|d|<0.5$ as small, $0.5\leq|d|<0.8$ as medium, and $|d|\geq 0.8$ as large.
Table I summarizes the statistical comparison between open and closed issues across all investigated features. All features show statistically significant differences ($p<0.01$), indicating that each characteristic varies meaningfully between open and closed issues. However, the effect size values ($d$) reveal that the magnitude of these differences is generally small or negligible, suggesting that while the differences are statistically detectable, their practical impact is limited.
However, the presence of labels and assigned developers (i.e., has-label and has-assignee) show small but meaningful effects ($d$ = 0.20 and 0.23, respectively). This finding aligns with expectations that labeled and assigned issues are more likely to attract attention and thus reach closure. These two features appear to play a modest but tangible role in issue management effectiveness. In contrast, textual and contextual attributes, such as title-length, body-length, has-code, and has-URLs exhibit negligible effect sizes ($d$ << 0.2), implying that the amount of written content or inclusion of supporting information (e.g., code snippets, URLs) does not substantially differentiate open from closed issues. Similarly, the number of contributors in a repository shows only a negligible effect, indicating that having more contributors does not necessarily lead to faster or more consistent issue resolution.
@@
Our analysis shows that management-related factors, particularly labeling and assignment, have the strongest association with issue closure in SWS repositories, albeit with small effect sizes. In contrast, content-related attributes (e.g., title length, body length, code snippets, URLs) and repository-level characteristics (e.g., number of contributors) show negligible influence.